# Rethinking Recurrent Neural Networks for Time Series Forecasting: A Reinforced Recurrent Encoder with Prediction-Oriented Proximal Policy Optimization

### v1

@@ Abstract
[label3]organization=School of Computer Science and Techonolgy, Huazhong University of Science and Technology, city=Wuhan, postcode=430074, state=Hubei, country=China
@@ 1 Introduction
Despite these advances, encoder-only RNN predictors remain limited in capturing the dynamic and diverse nature of temporal dependencies. As shown in Fig. 2(a), when future observations exhibit peak patterns (orange line), effective forecasting requires adaptively extracting salient characteristics from pattern similar segments in the historical inputs (highlighted by two orange dashed boxes). A shifted example from the same dataset ( Fig. 2(b)) further illustrates that different prediction segments rely on distinct historical patterns (highlighted by three green dashed boxes). These contrasting instances highlight the need for adaptive retrieval of relevant historical inputs and dynamic alignment between historical contexts and future targets. However, conventional encoder only RNNs, with their fixed recurrent transitions, lack flexible adjustment mechanisms, which constrains their ability to respond to evolving temporal structures [Oreshkin et al., 2020]. Therefore, developing adaptive learning mechanisms that enables dynamic alignment to accommodate time-varying patterns is crucial to advancing time series forecasting, which constitutes the focus of this study.
@@
The rest of this paper is organized as follows. Section 2 introduces the problem formulation and the literature related to this work. In Section 3, the proposed method is illustrated in detail. Section 4 introduces the data resources and experimental setup. Section 5 analyses the experimental results. Finally, the conclusion is drawn in Section 6.
@@ 2.2 Dynamic input feature selection
Heuristic methods formulate input selection as a combinatorial optimization problem aimed at identifying the optimal feature subset. Swarm intelligence algorithms, such as Particle Swarm Optimization (PSO) and its Boolean variants, have been integrated with deep learning models to improve generalization performance [Tijjani et al., 2024]. Genetic Algorithms (GA) is employed by using encoding, crossover, and mutation to efficiently explore high-dimensional feature spaces [Ileberi et al., 2022], while Whale Optimization Algorithm (WOA) is utilized by deriving sparse yet robust feature subsets [Nadimi-Shahraki et al., 2022]. However, these methods produce deterministic feature subsets prior to model training, limiting their adaptability to time-varying feature relevance. Moreover, their search complexity increases exponentially with feature dimensionality, and their capacity to capture cross-temporal dependencies in non-stationary time series remains limited [Xue et al., 2015].
@@ 2.3 Hidden feature learning mechanism
Compared to the extensive studies on hidden state skip connections, output target selection has received limited attention from researchers. In RNN-based forecasting models, output target selection determines the supervised relationships between hidden representations and prediction targets, thereby playing a crucial role in aligning learned temporal features with forecasting objectives. Traditional RNNs employ fixed linear transformations from final hidden states, which may lead to suboptimal modeling when the representations contain redundant information. To enhance adaptive hidden-output mapping, Krueger et al. [2017] introduce Zoneout, which stochastically preserves hidden activations. Wang et al. [2019] apply backtracking search to optimize output weights, while Sima et al. [2025] explore out mappings by combing the PSO algorithm with Bayes optimization. Nevertheless, existing work treats output mapping in isolation, neglecting its synergy with input selection and hidden-state connectivity. Therefore, this study aims to bridge this gap by coordinating adaptation across all three components, which is essential for capturing dynamic patterns in time series forecasting.
@@ 3.2.2 Dynamic transition sampling
Let $g\in\{1,\ldots,G_{\pi}\}$ denote the current epoch for training the agent, where $G_{\pi}$ is the total number of epochs. To further prevent premature convergence and enable periodic exploration resurgence, we incorporate cyclical annealing scheme that modulates temperatures over the course of training:
@@ 4.4 Implementation details
In all experiments, each dataset is scaled to [1,1][-1,1] by the min-max method, and is split into training set, validation set, and testing set with a ratio of 7:1:2. All methods are trained on the training set using the Adam optimizer, with the maximum number of training epochs set to 50 for agent networks and 20 for the RNN environment network. An early stopping mechanism is employed through cross-validation on validation set with a patience of 6 epochs to save the best performing RNN model across all methods. During evaluation on the testing set, predictions are transformed back to the original scale and then evaluated using the MSE and MAE metrics. All experiments are run 10 times with distinct random seeds on an NVIDIA RTX A4500 20GB GPU, and the mean performance across runs is reported.
@@ 5.2 Ablation study
To justify the individual contributions of the proposed RRE framework and PPO4Pred algorithm, we conduct an ablation study with with the variants described in Table 8. We design TA-PSO as a heuristic-based variant that employs the PSO algorithm [Kennedy and Eberhart, 1995] to optimize our designed Ternary Action (TA) from Section 3.1.1. Additionally, we develop three RL-based ablative variants, RRE-PG, RRE-DQN, and RRE-PPO, which retain the RRE framework but replace the PPO4Pred algorithm with alternative RL algorithms. RRE-PG uses vanilla Policy Gradient [Sutton et al., 2000], a foundational policy-based RL method without value function estimation. RRE-DQN uses Deep Q-Network [Mnih et al., 2015], a value-based RL method that learns action-value functions. RRE-PPO uses standard PPO [Schulman et al., 2017] without our prediction-oriented modifications. These RL-based variants isolate the contribution of our PPO4Pred algorithm by maintaining the same action space and framework while varying only the optimization method.
@@
    Among the three adopted RL algorithms, RRE-PG shows inferior performance compared to RRE-DQN and RRE-PPO, while RRE-DQN achieves comparable accuracy to RRE-PPO in most cases. This performance hierarchy demonstrates that the complexity of our optimization problem requires sophisticated RL algorithms with advanced modifications, justify the necessity of employing a strong RL algorithm within our designed RRE framework, rather than relying on basic policy-based methods.
@@ 5.3 Hyperparameter sensitivity and convergence analysis
The results in Fig. 11 demonstrate demonstrate clear convergence patterns across all five datasets. Both training and validation losses exhibit consistent downward trends across all datasets, validating the effectiveness of our asynchronous optimization strategy. Most datasets achieve convergence within 8–10 rounds, with the loss curves gradually stabilizing after approximately 10–15 rounds, suggesting that the architecture-predictor coevolution has reached near-optimal configurations. Notably, the observed divergence between training and validation curves in later rounds (typically after round 10–12) indicates the onset of overfitting, highlighting both the necessity and effectiveness of our early stopping mechanism. By continuously monitoring validation performance, our method can automatically terminate training at the point where generalization capability peaks depending on dataset characteristics.
@@
5.4 Computation efficiency analysis
@@
Comapred to LSTMjump and PGLSTM, our method incurs a moderately higher training time due to the asynchronous update mechanism, which alternates between predictor and agent update in an asynchronous manner, whereas both LSTMjump and PGLSTM update the predictor and agent synchronously. Additionally, our method has a larger parameter overhead compared to LSTMjump and PGLSTM. This is due to our Transformer-based PPO agent, which includes multi-head attention mechanisms and feed-forward layers, whereas LSTMjump and PGLSTM use simple MLP-based agent.
@@
We further visualizes the trade-off between the accuracy gain and computational cost in Fig. 12. Compared to the baseline methods, our RRE-PPO4Pred method consistently achieves lower MSE. As there is no free lunch in model optimization, the improved accuracy of our method comes at the cost of increased training time and parameters. However, as modern GPUs typically have 6–24 GB of memory, this parameter overhead, approximately 0.82 GB GPU memory for training and 0.48 GB for inference [^9] All GPU memory measurements are recoded with PyTorch v2.5.0 and CUDA v12.1 on NVIDIA RTX A4500 GPU., can be acceptable in many practical applications. Moreover, our training time remains significantly lower than that of EBPSO, and our T/P ratio is comparable to Naive-all and Naive-last baselines, demonstrating a favorable balance between accuracy and efficiency.
@@ 5.5 Performance comparison with Transformer-based methods
While the previous experiments focused on comparisons with RNN-based methods and ablation studies, we further examines whether our proposed RRE-PPO4Pred method can enable RNN-based models to compete with or outperform modern Transformer architectures, which representing recent advances in time series forecasting.
Specifically, Specifically, we include two representative Transformer-based methods, PatchTST [Nie et al., 2022] and iTransformer [Liu et al., 2023]. PatchTST applies patching and channel-independence mechanisms to achieve state-of-the-art performance on various benchmarks, and iTransformer [Liu et al., 2023] modifies the traditional Transformer-based forecasting model by applying attention across variates rather than time steps. Thus, we perform the experiments on the multivariate dataset ETTh, and multiple forecasting horizons are selected to evaluate performance across both short-term and long-term prediction scenarios. According to previous results in Table 5, we instantiate our RRE-PPO4Pred with xLSTM backbone, which shows an moderate performance among the selected recurrent architectures, avoiding cherry-picking the best-performing backbone and providing a more conservative evaluation of our method’s effectiveness. For further comparison, we include a vanilla xLSTM model optimized using the Naive-all baseline, which represents the standard xLSTM without policy-based modifications.
@@ 5.6 Visualization analysis
Across all subplots in Fig. 13, our method substantially reduce the occurrence of high-error instances, and consistently exhibits the most concentrated error distribution with the lowest error expectation. This distribution indicates fewer catastrophic prediction failures, which is crucial for reliable deployment in real-world applications. These findings demonstrate that our method achieves superior robustness and generalization capability compared to strong baseline methods, effectively addressing the challenge of instance-level variations in time series forecasting.

### v2

@@ Abstract
[label3]organization=School of Computer Science and Technology, Huazhong University of Science and Technology, city=Wuhan, postcode=430074, state=Hubei, country=China
@@ 1 Introduction
Despite these advances, encoder-only RNN predictors remain limited in capturing the dynamic and diverse nature of temporal dependencies. As shown in Fig. 2(a), when future observations exhibit peak patterns (orange line), effective forecasting requires adaptively extracting salient characteristics from pattern similar segments in the historical inputs (highlighted by two orange dashed boxes). A shifted example from the same dataset ( Fig. 2(b)) further illustrates that different prediction segments rely on distinct historical patterns (highlighted by three green dashed boxes). These contrasting instances highlight the need for adaptive retrieval of relevant historical inputs and dynamic alignment between historical contexts and future targets. However, conventional encoder only RNNs, with their fixed recurrent transitions, lack flexible adjustment mechanisms, which constrains their ability to respond to evolving temporal structures [Oreshkin et al., 2020]. Therefore, developing adaptive learning mechanisms that enable dynamic alignment to accommodate time-varying patterns is crucial for advancing time series forecasting, which constitutes the focus of this study.
@@
The rest of this paper is organized as follows. Section 2 introduces the problem formulation and the literature related to this work. In Section 3, the proposed method is illustrated in detail. Section 4 introduces the data resources and experimental setup. Section 5 analyzes the experimental results. Finally, the conclusion is drawn in Section 6.
@@ 2.2 Dynamic input feature selection
Heuristic methods formulate input selection as a combinatorial optimization problem aimed at identifying the optimal feature subset. Swarm intelligence algorithms, such as Particle Swarm Optimization (PSO) and its Boolean variants, have been integrated with deep learning models to improve generalization performance [Tijjani et al., 2024]. Genetic Algorithm (GA) is employed by using encoding, crossover, and mutation to efficiently explore high-dimensional feature spaces [Ileberi et al., 2022], while Whale Optimization Algorithm (WOA) is utilized by deriving sparse yet robust feature subsets [Nadimi-Shahraki et al., 2022]. However, these methods produce deterministic feature subsets prior to model training, limiting their adaptability to time-varying feature relevance. Moreover, their search complexity increases exponentially with feature dimensionality, and their capacity to capture cross-temporal dependencies in non-stationary time series remains limited [Xue et al., 2015].
@@ 2.3 Hidden feature learning mechanism
Compared to the extensive studies on hidden state skip connections, output target selection has received limited attention from researchers. In RNN-based forecasting models, output target selection determines the supervised relationships between hidden representations and prediction targets, thereby playing a crucial role in aligning learned temporal features with forecasting objectives. Traditional RNNs employ fixed linear transformations from final hidden states, which may lead to suboptimal modeling when the representations contain redundant information. To enhance adaptive hidden-output mapping, Krueger et al. [2017] introduce Zoneout, which stochastically preserves hidden activations. Wang et al. [2019] apply backtracking search to optimize output weights, while Sima et al. [2025] explore output mappings by combining the PSO algorithm with Bayes optimization. Nevertheless, existing work treats output mapping in isolation, neglecting its synergy with input selection and hidden-state connectivity. Therefore, this study aims to bridge this gap by coordinating adaptation across all three components, which is essential for capturing dynamic patterns in time series forecasting.
@@ 3.2.2 Dynamic transition sampling
Let $g\in\{1,\ldots,G_{\pi}\}$ denote the current epoch for training the agent, where $G_{\pi}$ is the total number of epochs. To further prevent premature convergence and enable periodic exploration resurgence, we incorporate a cyclical annealing scheme that modulates temperatures over the course of training:
@@ 4.4 Implementation details
In all experiments, each dataset is scaled to [1,1][-1,1] by the min-max method, and is split into training set, validation set, and testing set with a ratio of 7:1:2. All methods are trained on the training set using the Adam optimizer, with the maximum number of training epochs set to 50 for agent networks and 20 for the RNN environment network. An early stopping mechanism is employed through cross-validation on the validation set with a patience of 6 epochs to save the best performing RNN model across all methods. During evaluation on the testing set, predictions are transformed back to the original scale and then evaluated using the MSE and MAE metrics. All experiments are run 10 times with distinct random seeds on an NVIDIA RTX A4500 20GB GPU, and the mean performance across runs is reported.
@@ 5.2 Ablation study
To justify the individual contributions of the proposed RRE framework and PPO4Pred algorithm, we conduct an ablation study with the variants described in Table 8. We design TA-PSO as a heuristic-based variant that employs the PSO algorithm [Kennedy and Eberhart, 1995] to optimize our designed Ternary Action (TA) from Section 3.1.1. Additionally, we develop three RL-based ablative variants, RRE-PG, RRE-DQN, and RRE-PPO, which retain the RRE framework but replace the PPO4Pred algorithm with alternative RL algorithms. RRE-PG uses vanilla Policy Gradient [Sutton et al., 2000], a foundational policy-based RL method without value function estimation. RRE-DQN uses Deep Q-Network [Mnih et al., 2015], a value-based RL method that learns action-value functions. RRE-PPO uses standard PPO [Schulman et al., 2017] without our prediction-oriented modifications. These RL-based variants isolate the contribution of our PPO4Pred algorithm by maintaining the same action space and framework while varying only the optimization method.
@@
    Among the three adopted RL algorithms, RRE-PG shows inferior performance compared to RRE-DQN and RRE-PPO, while RRE-DQN achieves comparable accuracy to RRE-PPO in most cases. This performance hierarchy demonstrates that the complexity of our optimization problem requires sophisticated RL algorithms with advanced modifications, which justifies the necessity of employing a strong RL algorithm within our designed RRE framework, rather than relying on basic policy-based methods.
@@ 5.3 Hyperparameter sensitivity and convergence analysis
The results in Fig. 11 demonstrate clear convergence patterns across all five datasets. Both training and validation losses exhibit consistent downward trends across all datasets, validating the effectiveness of our asynchronous optimization strategy. Most datasets achieve convergence within 8–10 rounds, with the loss curves gradually stabilizing after approximately 10–15 rounds, suggesting that the architecture-predictor coevolution has reached near-optimal configurations. Notably, the observed divergence between training and validation curves in later rounds (typically after round 10–12) indicates the onset of overfitting, highlighting both the necessity and effectiveness of our early stopping mechanism. By continuously monitoring validation performance, our method can automatically terminate training at the point where generalization capability peaks depending on dataset characteristics.
@@
5.4 Computational efficiency analysis
@@
Compared to LSTMjump and PGLSTM, our method incurs a moderately higher training time due to the asynchronous update mechanism, which alternates between predictor and agent update in an asynchronous manner, whereas both LSTMjump and PGLSTM update the predictor and agent synchronously. Additionally, our method has a larger parameter overhead compared to LSTMjump and PGLSTM. This is due to our Transformer-based PPO agent, which includes multi-head attention mechanisms and feed-forward layers, whereas LSTMjump and PGLSTM use simple MLP-based agent.
@@
We further visualize the trade-off between the accuracy gain and computational cost in Fig. 12. Compared to the baseline methods, our RRE-PPO4Pred method consistently achieves lower MSE. As there is no free lunch in model optimization, the improved accuracy of our method comes at the cost of increased training time and parameters. However, as modern GPUs typically have 6–24 GB of memory, this parameter overhead, approximately 0.82 GB GPU memory for training and 0.48 GB for inference [^9] All GPU memory measurements are recorded with PyTorch v2.5.0 and CUDA v12.1 on NVIDIA RTX A4500 GPU., can be acceptable in many practical applications. Moreover, our training time remains significantly lower than that of EBPSO, and our T/P ratio is comparable to Naive-all and Naive-last baselines, demonstrating a favorable balance between accuracy and efficiency.
@@ 5.5 Performance comparison with Transformer-based methods
While the previous experiments focused on comparisons with RNN-based methods and ablation studies, we further examine whether our proposed RRE-PPO4Pred method can enable RNN-based models to compete with or outperform modern Transformer architectures, which represent recent advances in time series forecasting.
Specifically, we include two representative Transformer-based methods, PatchTST [Nie et al., 2022] and iTransformer [Liu et al., 2023]. PatchTST applies patching and channel-independence mechanisms to achieve state-of-the-art performance on various benchmarks, and iTransformer [Liu et al., 2023] modifies the traditional Transformer-based forecasting model by applying attention across variates rather than time steps. Thus, we perform the experiments on the multivariate dataset ETTh, and multiple forecasting horizons are selected to evaluate performance across both short-term and long-term prediction scenarios. According to previous results in Table 5, we instantiate our RRE-PPO4Pred with xLSTM backbone, which shows a moderate performance among the selected recurrent architectures, avoiding cherry-picking the best-performing backbone and providing a more conservative evaluation of our method’s effectiveness. For further comparison, we include a vanilla xLSTM model optimized using the Naive-all baseline, which represents the standard xLSTM without policy-based modifications.
@@ 5.6 Visualization analysis
Across all subplots in Fig. 13, our method substantially reduces the occurrence of high-error instances, and consistently exhibits the most concentrated error distribution with the lowest error expectation. This distribution indicates fewer catastrophic prediction failures, which is crucial for reliable deployment in real-world applications. These findings demonstrate that our method achieves superior robustness and generalization capability compared to strong baseline methods, effectively addressing the challenge of instance-level variations in time series forecasting.
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 71931005, the China Postdoctoral Science Foundation under Grant 2024M761027, and the Hubei Provincial Natural Science Foundation of China under Grant 2025AFB110.